Develop a lot-for-lot solution and calculate total relevant costs for the gross requirements in the following table*. PeriodB.   TOTAL COST________
Answer #1

Answers are in BOLD

Period 1 2 3 4 5 6 7 8 9 10 11 12 Total cost
Gross requirements GR 30 0 30 0 30 60 20 0 20 70 0 60
Scheduled receipts SR 0 0 0 0 0 0 0 0 0 0 0 0
safety stock, SS=0 on hand at the end PIOH 10 10 0 0 0 0 0 0 0 0 0 0
Beginning inventory-280 on hand at the beginning IOH 40 10 10 0 0 0 0 0 0 0 0 0
Requirement -10 -10 20 0 30 60 20 0 20 70 0 60
Net requirements NR 0 0 20 0 30 60 20 0 20 70 0 60
Q = 100 Planned order receipts (PR) 0 0 20 0 30 60 20 0 20 70 0 60
Lead Time LT=1 Planned order releases POR 0 20 0 30 60 20 0 20 70 0 60 0
Holding cost= $1.5* PIOH 15 15 0 0 0 0 0 0 0 0 0 0 30
setup cost (whenever there is new order reciept) 0 0 200 0 200 200 200 0 200 200 0 200 1400

Answer b: Total cost= cost of setup + cost of holding imventory= 30+1400= $1430

repeat these steps for every week
step 1 IOH= PIOH of last week
step 2 Requirement= GR-IOH-SR+SS
step 3 NR= is equal to requirement (if requirement>0) OR 0 (if requirement is negative)
step 4 PR= in L4L PR=NR more than or equal to NR in multiples of lot size
step 5 PIOH= IOH+PR-GR-SR
step 6 POR= PR of preceding week
step 7 (After completing whole MRP) Holding cost= 1.5$*PIOH

Please ask, if you have any doubts through the comment section. Do rate the answer